Sales $ 397,596 $ 261,576 $ 208,427 $ 149,410 $ 111,500
Cost of goods sold 209,250 137,520 111,682 79,143 57,980
Accounts receivable 19,283 15,302 14,256 8,711 7,627


Compute trend percents for the above accounts, using 2015 as the base year.

Trend Percent for Net Sales:
Choose Numerator: / Choose Denominator:
/ = Trend percent
2019: / = %
2018: / = %
2017: / = %
2016: / = %
Trend Percent for Cost of Goods Sold:
Choose Numerator: / Choose Denominator:
/ = Trend percent
2019: / = %
2018: / = %
2017: / = %
2016: / = %
Trend Percent for Accounts Receivables:
Choose Numerator: / Choose Denominator:
/ = Trend percent
2019: / = %
2018: / = %
2017: / = %
2016: / = %
Answer #1
Trend Percent for Net Sales:
Choose Numerator: / Choose Denominator:
Current year sales / Base year sales = Trend percent
2019: 397,596 / 111,500 = 356.6%
2018: 261,576 / 111,500 = 234.6%
2017: 208,427 / 111,500 = 186.9%
2016: 149,410 / 111,500 = 134%
Trend Percent for Cost of Goods Sold:
Choose Numerator: / Choose Denominator:
Current year Cost of Goods Sold / Base year Cost of Goods Sold = Trend percent
2019: 209,250 / 57,980 = 360.9%
2018:
137,520
/ 57,980 = 237.2%
2017: 111,682 / 57,980 = 192.6%
2016: 79,143 / 57,980 = 136.5%
Trend Percent for Accounts Receivables:
Choose Numerator: / Choose Denominator:
Current year accounts receivables / Base year accounts receivables = Trend percent
2019: 19,283 / 7,627 = 252.8%
2018: 15,302 / 7,627 = 200.6%
2017: 14,256 / 7,627 = 186.9%
2016: 8,711 / 7,627 = 114.2%

Exact answer may slightly differ due to rounding off.

There was no instruction regarding rounding off, so I rounded off all figures upto 1 decimal place.

Kindly give a positive rating if you are satisfied with the answer. Feel free to ask if you have any doubt. Thanks